What is the magnetic field strength (in T) inside an ideal solenoid with 2650 loops and a length of 35 cm that is carrying a current of 0.385 A? Answer:

please show steps and units.
Answer #1

magnetic field B = u0*(N/L)*I

B = 4*(22/7)*10^(-7)*(2650/0.35)*0.385 =0.00366 T

B = 0.00366 T = 3.66*10^(-3) T